1/1

William Wordsworth,  William Wordsworth,

Evening on Calais Beach Ein Abend an der Küste von Calais

tradukita de Manfred Retzlaff

It is a beauteous evening, calm and free, The holy time is quiet as a Nun Breathless with adoration; the broad sun Is sinking down in its tranquillity; Ein schöner Abend, Ruhe um mich her. Anbetend still verharrt die heil’ge Zeit, Kein Lufthauch ist zu spüren weit und breit, Rot glühend sinkt die Sonne in das Meer.

The gentleness of heaven broods o’er the sea: Listen! the mighty Being is awake, And doth with his eternal motion make A sound like thunder—everlastingly.

Des Himmels Güte lastet milde dort.

Doch horch! Das mächt’ge Wesen ist erwacht Und lässt mit seiner ew’gen Himmelsmacht Ein fernes Donnern hören - immerfort.

Dear Child! dear Girl! that walkest with me here, If thou appear untouch’d by solemn thought, Thy nature is not therefore less divine:

................ ................ ................

Thou liest in Abraham’s bosom all the year; And worshipp’st at the Temple’s inner shrine, God being with thee when we know it not.

................ ................ ................

Verkinto de tiu ĉi Angla poemo estas William Wordsworth (∗1770 – †1850).

Arg-312-630 (2005-05-19 10:27:47)

Traduko de la Angla poemo “Evening on Calais Beach” de William Wordsworth (∗1770 – †1850) en la Germanan de Manfred Retzlaff (∗1938-11-04) en 2005-05-19.

Arg-312-631 (2005-05-19 10:24:37)

http://www.poezio.net/version?poem-id=312&version-id=630,631